Project

General

Profile

2015-09-03-MIPO » History » Version 1

Whyte, Colin, 03 September 2015 19:29

1 1 Whyte, Colin
Present CW, SW, MP, PG, RK, JP, PS
2
 
3
Agenda
4
1.	MEMO report – no meeting to report, next MEMO scheduled 15/09/15
5
2.	Actions
6
a.	Offset Bellows. Discussed design with AN and JT, prefer concentric bellows assembly with slotted bolt holes. Contacted Hyspan, thanks to Steve Virostek for details, no reply as yet. Also contacted suppliers in UK. At least 1 possible. Update post meeting: delivery longer than expected ~10-12 weeks – CW.
7
b.	Action on SW to estimate times – completed: in last meeting’s minutes
8
c.	Action on CW to clarify BOB button via Craig MacWaters – completed Bob Buttons do drop SC permit 
9
d.	Action on CW to contact TH – completed.
10
e.	Action on PMH logging sheets to go to experts this week, ongoing
11
f.	Action on CW to photograph and circulate access platform, superceeded. 
12
g.	Action on CW to discuss LiH secondary abs with RP - completed.
13
14
3.	FC status
15
Valve for stinger port arrived today - MT fitting now. 
16
02/09/2015 JC and MC made RGA reading of insulating vacuum space
17
03/09/2015 N2 purge of insulating vac. Repeat RGA measurement 04/09/2015.
18
Unlikely to cool 04/09/2015. – FC moving on to critical path. FC on 3.5 week readiness trajectory, assuming successful cooldown.
19
AB arrives 04/09/2015 will consult.
20
Contact established with Tesla, a representative will be on site 04/09 on another matter SW hopes to consult. SW and SH to visit Tesla next week
21
4.	SS training
22
SSD, 212A 5 quenches. On 6th cycle there was a problem, magnet was ramped down resulting in a quench (not considered a training quench)
23
SSU, repaired, warm check out, cooled down, fully cold weekend 22/08
24
He line crack at weld - replaced. 
25
Training last week, made 6 quenches.
26
5.	Magnet electrical
27
Power supply issues, original diagnosis was quench system. Problem appears to be related to resistance in high current conductors between power supplies and magnets. System moves outside specifications over 175-180A. 
28
Various patches under evaluation by DL and FNAL staff. 
29
MP/AB should be included in all magnet discussions. 
30
MP requires a full conceptual review on integrated magnets and power supplies. 
31
Some options may have significant implementation times. 
32
MP meeting Tues Afternoon with DL/SG face to face at RAL. Magnet experts also required.
33
Action on CW to arrange meeting at RAL Tues 8th September. 
34
35
To proceed must reduce ramp rate to 12mA/s, ramp will require 6-7 hrs. 
36
Training would require ramp from 6am, afternoon access.  Hall may have to close again by 3pm for next quench.
37
On this basis AB and DB will arrive tomorrow. 
38
39
PS is it possible to run for low momentum particles? JP for 140MeV/C particles we require 198A. MP current match coil maximum current is 180A
40
MP should not try to run at low currents as this could delay solution. JP agrees, important UK review in Jan/Feb. MP important Mice MPB mid Oct 2015.
41
.
42
43
6.	End of construction? N/A 
44
7.	Activation study 29th August – not completed ISIS off. 
45
8.	RF Review.
46
To be held at Cosners House.
47
Critical review to examine construction, integration and commissioning. 
48
Review report will feed into MPB in Oct 2015 
49
9.	AOB.